how do you solve
x+3=23
3x-2y=3
: how do you solve
x+3=23
3x-2y=3

Answer by checkley77(3848)

x+3=23 or x=23-3 or x=20 ans.
3x-2y=3
3*20-2y=3
60-2y=3
-2y=3-60
-2y=-57
-y=-57/2
y=57/2 ans.